HanAkoSolo (jamierox4ev3r):

Math Review Day 3! 6e. Rationalize the expression and simplify.

HanAkoSolo (jamierox4ev3r):

Let me re-type the question 6e. Rationalize the expression and simplify \(\huge\frac{\sqrt{4+h}-2}{h}\)

OpenStudy (anonymous):

Is this root just for 4? Or h also?

HanAkoSolo (jamierox4ev3r):

the root is \(\sqrt{4+h}\) @joyraheb

Nnesha (nnesha):

i guess multiply the numerator and the denominator by the `conjugate` of the numerator

HanAkoSolo (jamierox4ev3r):

I do believe that the first step would be to multiply the whole expression by the conjugate of the numerator. So I assume that it would look a little like this: \(\huge\frac{\sqrt{4+h}-2}{h}\times\huge\frac{\sqrt{4+h}+2}{\sqrt{4+h}+2}\)

HanAkoSolo (jamierox4ev3r):

LOL you read my mind XD There is a lot of multiplying of conjugates in Algebra. A common trend that I recall

HanAkoSolo (jamierox4ev3r):

If you multiply these, you should get \(\huge\frac{(4+h)-4}{h(\sqrt{4+h}+2}\) right?

Nnesha (nnesha):

yes right!

HanAkoSolo (jamierox4ev3r):

yes! oops wait i think a parentheses is missing in in the denominator \(\huge\frac{(4+h)-4}{h(\sqrt{4+h})+2}\)

HanAkoSolo (jamierox4ev3r):

anyhow, from here I'm a little stuck. I know the in the numerator, 4 and -4 cancel out, so you're left with h...but what happens in the denominator?

Nnesha (nnesha):

yes right \(\huge\rm \frac{\color{reD}{4}+h\color{reD}{-4}}{h(\sqrt{4+h}+2)}\)\[\rm \frac{ h }{ h(\sqrt{4+h}+2)}\] can you cancel out anything ?

HanAkoSolo (jamierox4ev3r):

oh wait! i see. So i mentioned canceling out in the numerator, but it seems like something in the numerator and denominator can be cancelled out \(\huge\rm \frac{\color{reD}{4}+\color{green}h\color{reD}{-4}}{\color{green}h(\sqrt{4+h}+2)}\) The things in green become 1, so wouldn't you have this: \(\huge\frac{1}{\sqrt{4+h}+2}\)

Nnesha (nnesha):

parentheses should be aftr 2 cuz it's one expression (sqrt{4+h}+2)

Nnesha (nnesha):

that's right!

HanAkoSolo (jamierox4ev3r):

oh true, whoops XD and ty ;-;

Nnesha (nnesha):

np don't cry:(

HanAkoSolo (jamierox4ev3r):

yeah i put the parentheses in the wrong place earlier. my bad hehe I'm fine. From there, am I done?

HanAkoSolo (jamierox4ev3r):

is that the final answer?

Nnesha (nnesha):

btw do you have the answer ? just want to know cuz most of times we r not allowed to leave the square root at the denominator

Nnesha (nnesha):

yes i think so we can't do anything else with that exprssion

HanAkoSolo (jamierox4ev3r):

hmmm lemme check online. The textbook doesn't have the answers in it, but the textbook's website has them. Brb

Nnesha (nnesha):

oh cool

HanAkoSolo (jamierox4ev3r):

Yep that's it. Thank you thank you! XD

Nnesha (nnesha):

my pleasure